OpenStudy (anonymous):

Simplify completely 6x-12/10

OpenStudy (anonymous):

Well, I don't see how we could answer this, (not an equation), apart from dividing the entire thing to get x. In this equation, it would be divided by 6.

OpenStudy (anonymous):

6x -12/10 = 2(3x-6)/10 = 3x-6/5

OpenStudy (anonymous):

Ah, let's take a bite of that. Firstly, we would need to simplify the middle part, which would involve multiplying 2 by 3x and 2 by negative 6, yielding \[(6x-12)/10\] This leaves us with \[(6x-12)/10=3x-6/5\] Simplify by dividing each part of the parenthesis by 10. \[6/10x-12/10=3x-6/5\]

OpenStudy (anonymous):

Subtract 3/5x and 6/5 from each side. This leaves us with \[9/5x=0\] Which doesn't make sense. Sorry for not being able to resolve it.

OpenStudy (anonymous):

it's okay im confused too

OpenStudy (anonymous):

is this right? 3(x-2)/5

OpenStudy (anonymous):

Anyone wana help me pleasse!